[Halla_running] [EXTERNAL] Hall A RC Daily Summary August 20th, 2025
McMurtry, Jacob Thomas (rby2vw)
rby2vw at virginia.edu
Wed Aug 20 08:45:15 EDT 2025
Dear Collaborators
Yesterday day shift was a rocky time for the accelerator. Last minute restricted access for hall C and B for dry runs of the Energy Secretary tour tomorrow were followed up a cryo trip which took beam down for most of the day shift (https://ace.jlab.org/dtm/all-events?qualified=true&incidentId=16138 ), followed by some scattered downtimes during the first half of swing shift. Beam initially came back around 5 but was spotty, shift crew attempted to do cross check during this time and were able to complete it around 7, beam finally stabilized around 9 and ran smooth through the OWL shift. During this time, we continued taking production data on LH2 with 15 uA beam. During late swing/early OWL, Vidura continued work on the GEM gas mixture (https://logbooks.jlab.org/entry/4446454 ), and yesterday day Jhih-Ying changed come hot channels on ECal (https://logbooks.jlab.org/entry/4445944 ). The current run plan is to continue taking production data at 15 uA on LH2.
Today is RF recovery day, MCC will not be opening up any of the injector, north & south linacs, or or BSY. They plan to be done by around 2:30 this afternoon, and are hoping to deliver beam sometime mid swing (7:30 or so). Halls B and D have requested rapid and controlled access for this time, hall C will be moving to restricted access. Hall A is scheduled to be surveyed at 10:00 this morning for controlled access, with the intent to lock up around 2:00.
Since yesterday's Day to today's owl:
* Total Accumulated Charge at Kin3 so far: 87.61 C (Goal: 100 C)
* Total Accumulated Charge at Kin3 since yesterday : 0.66 C
Updates from MCC:
* Today is RF recovery day
* Compton test is scheduled for Thursday day shift (estimated 8 hours, no other beam studies planned), to begin around 8:30 and end around 16:30. There will likely be a break around 14:10 due to the Energy Secretary tour at MCC
* The Energy Secretary tour is this Thursday, hall C will be surveyed around 7:30 for the tour at 10:40, hall B surveyed at 11:30 for tour at 14:35, MCC tour at 14:10.
Current Issues and Hall activities:
* Ongoing GEM gas tests
* Target leak/pump failure
Shift Summary:
* Yesterday's Day: https://logbooks.jlab.org/entry/4445883
* Yesterday's Swing: https://logbooks.jlab.org/entry/4446228
* Today's Owl: https://logbooks.jlab.org/entry/4446419
RC Meeting:
* We will have our usual RC meeting today at 1 pm in the conference room in the second floor of the CH and on Zoom.
* https://urldefense.proofpoint.com/v2/url?u=https-3A__jlab-2Dorg.zoomgov.com_j_1616220064&d=DwIFAw&c=CJqEzB1piLOyyvZjb8YUQw&r=cE_5H9I241Ivk1xpqB8Fq0UiLTkf4zMIQ6sedLYb2m8&m=dw2VrLVYeV8eD8nO59-ed98oNOORFZYHE3Kxqh39o_6nTPyfCsdFZHBcx7fO-xun&s=2MQHZi8VP-_TXMCJQEvCMfsmZsi8lq-wckSPfcqpRtc&e=
* Meeting ID: 161 622 0064
* Passcode: 314159
Best,
Jacob McMurtry
Hall A RC
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/halla_running/attachments/20250820/367702f0/attachment.htm>
More information about the Halla_running
mailing list